OpenStudy (anonymous):

-1/3(n+8) + 1/3n ≤ 1-n what is n ??

OpenStudy (anonymous):

\[\frac{ -1 }{3 }(n+8)+\frac{ 1 }{ 3 }n \le 1-n\] distribute \[\frac{ -1}{3} n + \frac{ 8}{ 3}+\frac{ 1}{ 3}n \le 1-n\] combine like terms \[\frac{ 8}{ 3 }\le1-n\] subtract 1 \[\frac{ 5}{ 3 }\le-n\] divide by -1 \[\frac{ -5}{ 3 }\le n\]